# Dissecting the SWE-Bench Leaderboards: Profiling Submitters and Architectures of LLM- and Agent-Based Repair Systems

### v1

@@ 3.2.1. Human-Workflow with Fixed execution - No Agent (G1)
The entries listed in this section follow a predefined workflow typically composed of two or three stages, such as localization, patch generation and validation. Moreover, none of these stages exhibit agentic behavior.
Jimenez et al. (jimenez2024SWEBenchLLMs, ), the creators of SWE-Bench, defined LLM-based approaches to try to solve the issues from the benchmark they had created. These entries are known as RAG + <model name>. The approach uses BM25 retrieval (robertson2009probabilistic, ) to retrieve candidate buggy files and then calls LLMs to generate patches. They also fine-tuned CodeLlama (roziere2024codellama, ) with 19,000 issues, giving as results the model SWE-LLAMA used for patch generation.
@@
Agentless (xia2024agentlessdemystifyingllmbasedsoftware, ) is a repair workflow based on LLMs, composed of three main stages: Localization, Repair and Validation (27.33% on Lite). This workflow is similar to traditional APR, such as those implemented in Astor framework (martinez2016Astor, ) or the state-of-the-art of the traditional APR TBar (Liu2019Tbar, ), with the novelty that uses LLMs to perform each task. As difference to LLM-based Agents, in Agentless, LLMs are not used to decide how to proceed in the repair workflow. Both leaderboards also includes submissions of the posterior version: Agentless-1.5 (40.67% on Lite, 50.8% on Verified).
@@
As a pioneering approach on SWE-Bench, Agentless has been extended, modified and improved, with the results of some of these variants also submitted to the leaderboards. For example, Agentless - Lite (32.3% on Lite, 42.4% on Verified), as the name suggest, is a simplification of the parent approach: “Use an embedding model to retrieve relevant files from the repository (Exclusively RAG-based localization). Query the LLM to generate a repair based on the top 5 retrieved files, retrying the generation until the model outputs a valid patch.”. The entry Agentless + Repograph (29.67% on Lite) corresponds to the evaluation Agentless augmented with the plugin Repograph. Repograph provides agents a structured way to analyze and interact with complex codebases, enabling detailed tracing of code dependencies, execution flow and structural relationships across the repository (ouyang2024repographenhancingaisoftware, ). Repograph information is integrated across all stages of the Agentless workflow.
@@
ORCALOCA (yu2025orcaLocallmAgent, ) is an approach focused on file and function localization, which is evaluated on top of Agentless. Instead of relying exclusively on LLMs to generate exploratory actions within the codebase, ORCALOCA introduces a dynamic action scheduling system that produces actions at varying levels of granularity and prioritizes them based on contextual relevance and urgency. The submission OrcaLoca + Agentless-1.5 (41% on Lite) incorporates ORCALOCA in the Localization stage from Agentless-1.5, while retaining the Repair, Patch Validation and Patch Selection stages. Similarly, CodeFuse-AAIS is other entry that relies on Agentless, and proposes two main improvements: 1) adaptive fault localization, and 2) expansion of the sampling space by combining multiples LLMs. . Nevertheless, we could not find more details about such improvements.
@@
SWE-Fixer (xie2025SWEFixerTrainingopensourcellms, ) (24.67% on Lite, 32.8% on Verified) is described by its authors as a further simplification of Agentless. Specifically, the approach follows a pipeline structure consisting of two primary stages: 1) code file retrieval, and 2) code editing. In both stages, it uses fine-tuned open-source models (Qwen 2.5) trained using Chain-of-Thought (CoT) data (wei2023chainofthought, ).
@@
Aider (26.33% on Lite) employs iterative process composed of the following main stages: 1) File Filtering:Aider builds a repository map, a compact and expressive summary of the entire codebase, using static code analysis. This map helps the LLM understand the project structure and dependencies. The model then produces a natural language response indicating which files need to be edited and why. 2) LLM-based Code Editing:Code edits are performed using a set of carefully designed prompting strategies and configurable editing backends. 3) Linting:A static analysis phase is used to ensure that the modified code has no syntax or other fatal errors. 4) Testing:The final step consists of running all available tests to verify whether the generated patches are plausible.
@@ 3.2.2. Human-Workflow with Fixed execution - with Single Agent (G2)
Meta introduced the SWE-RL approach (wei2025SWE-RL, ) to train LLMs —resulting in the model Llama3-SWE-RL— with the goal of enabling them to autonomously recover and replicate a developer’s reasoning process. To evaluate the ability of the trained models to fix issues, the authors generated a pipeline-based scaffold named Agentless Mini with include capabilities such as file-level fault localization, generating repair edits, reproduction, test generation, and regression test selection. Agentless Mini enhances the original Agentless’ pipeline with two key improvements: 1) test generation by considering other tests, and 2) patch selection based on voting. The evaluation of this scaffold using the trained model achieved a 41.2% on SWE-Bench Verified.
@@ 3.2.3. Human-Workflow with Fixed execution - with Multiple Agents (G3)
MASAI (arora2024masaimodulararchitecturesSE, ) (27.33,32.6% on Verified) presents a high-level modular architecture composed of five distinct stages, including Edit Localizer, Fixer and Ranker. Each stage is executed by a dedicated LLM-powered sub-agent, which is instantiated with defined objectives and tailored strategies designed to fulfill those objectives. In particular, each sub-agent is equipped with a set of actions and a corresponding strategy, a problem-solving approach for using the LLM to address its assigned sub-task. This strategy can include methods such as vanilla completion, Chain-of-Thought(wei2023chainofthought, ) or ReAct (yao2023ReAct, ). Note that in MASAI, agentic behavior is confined to individual stages, with each sub-agent operating independently within its assigned role, rather than exercising control over the overall system workflow.
@@ 3.2.5. Human-Workflow with Scaffolded Execution - Multiple Agents (G5)
Other approaches that report results on SWE-Bench using agentic workflows with multiple-agents, but with less publicly available detail, include Bracket.sh (53.2% on Verified),Composio SWE-Kit (41% on Lite, 48.6% on Verified).
@@ 3.2.6. Workflow Emergent and Emergent Autonomy - Single Agent (G6)
A similar approach, but using different LLMs, corresponds to the entries W&B Programmer O1 (64.6% on Verified), which was the first on SWE-Bench to evaluate the reasoning capabilities of OpenAI’s GPT-o1, and Google Jules (on Verified), which uses Google’s Gemini model.
@@ 3.2.8. Unclassified Submissions Due to Limited Evidence on Architectural Aspects (G8)
3.3. Architecture of solutions and % Resolved
@@
% Resolved % Resolved
Architecture #E Median Max #E Median Max
G1 Human-Workflow with Fixed execution - No Agent 20 26.83 41 13 26.2 50.8
G2 Human-Workflow with Fixed execution - with Single Agent - - - 4 49.7 68.2
G3 Human-Workflow with Fixed execution - with Multiple Agents 5 27.33 39.33 3 50 63.4
G4 Human-Workflow with Scaffolded Execution - Single Agent 14 36.84 48.67 7 55 65.4
G5 Human-Workflow with Scaffolded Execution - Multiple Agents 9 30.67 48.33 13 40.6 53.2
G6 Workflow Emergent and Emergent Autonomy - Single Agent 7 21.67 41.67 17 49 65.8
@@
G8 Unclassified Submissions 11 38.33 55 18 44.8 66.4
Table 9. Total (#E), Median and Max precision (Resolved) of submissions grouped by the architecture.
Table 9 groups the submission according to the classification of the architectures discussed previously. For each group, it shows the total number of submissions, and the median and max precision (Resolved). In SWE-Bench Lite the groups corresponding to the unclassified submissions (G8) hold the higher median and max precision (38 and 55 respectively). Human-Workflow with Scaffolded Execution - Single Agent (G4) holds the second place in all number of submissions, median and max precision. From the classified submission in Lite, architectures with scaffolded execution and agent involvement such as G4 and G5, achieve the best results. We assess whether there are statistically significant differences in precision across the groups. The Kruskal–Wallis test reveals a statistically significant difference between at least two groups of submitters (H = 19.4149, p = 0.0070). Dunn’s post-hoc test identifies a significant difference between G1 and G8 ($p\_value$ = 0.03473) and G6 and G8 ($p\_value$ = 0.04154)
In SWE-Bench Verified, two groups have the highest results, first (G4: Human-Workflow with Scaffolded Execution - Single Agent) with 55% followed byG7: Workflow Emergent and Emergent Autonomy - Multiple Agents he Kruskal–Wallis test reveals a statistically significant difference between at least two groups of submitters (H = 19.4149, p = 0.0070). Dunn’s post-hoc test identifies a significant difference between G1 and G8 ($p\_value$ = 0.04712) and G1 and G6 ($p\_value$ = 0.03289)
Answer to RQ 2 : On the SWE-Bench Lite leaderboard, the highest median precision is observed in the Unclassified Submissions (G8, 38.3%), followed by Human-Workflow with Scaffolded Execution – Single Agent (G4, 36.8%). In SWE-Bench Verified, Human-Workflow with Scaffolded Execution – Single Agent (G4) achieves the highest median precision (55%), followed by Workflow Emergent – Multiple Agents (G7, 54.5%). The largest groups in terms of submissions are G6 (Workflow Emergent and Emergent Autonomy - Single Agent, 18). Overall, scaffolded workflows —particularly with a single agent— consistently yield high precision across both leaderboards. Submissions without agents (G1) tend to show lower precision, but still contribute a substantial number of entries.
@@
We describe the end-to-end software maintenance pipeline proposed in the survey by Liu et al. (liu2024SurveyAgentsLLM4SE, ), which also classifies six agent-based systems that we examine in this paper: Agentless (xia2024agentlessdemystifyingllmbasedsoftware, )AutoCodeRover(Zhang2024AutoCoderRover, ), RepoUnderstander (subsequently renamed LingmaAgent(OLDma2024understandRepoUnderstander, ), MASAI (arora2024masaimodulararchitecturesSE, ), SWE-Agent(yang2024sweagentagent, ) and CodeR (xia2024dcoder, ).
@@ 3.4.1. Preprocessing
One of the initial phases commonly employed across approaches is the preprocessing of the codebase to construct a condensed representation, such as a graph or an index, that summarizes key information from the repository, including source code and, in some cases, documentation. This representation enables more efficient code navigation and supports contextual understanding.
@@
Knowledge graph representations of the codebase are employed by multiple approaches. In their survey, Liu et al. (liu2024SurveyAgentsLLM4SE, ) report one, RepoUnderstander, which builds a knowledge graph to facilitate software comprehension. We identified additional approaches evaluated on SWE-Bench that also adopt this strategy. For instance, the MarsCode Agent (liu2024marscode, ) analyzes and organizes the code and documentation within a repository to generate a multi-directional graph. Similarly, Nemotron-CORTEXA constructs a graph-based representation of a software repository, where nodes correspond to code files and entities —such as functions, classes, and methods— and edges capture contain and use relationships. Orcaloca (yu2025orcaLocallmAgent, ) employs a graph-based representation, called CodeGraph, to index code repositories and support code entity search. This graph captures hierarchical relationships —such as methods within classes— as well as reference relationships, including function calls between entities. Similarly, DARS (aggarwal2025darsdynamicactionresampling, ) utilizes RepoGraph (ouyang2024repographenhancingaisoftware, ), a hierarchical structure in which nodes represent code definitions and edges encode dependencies among them.
@@
Repository mapping is also adopted as a strategy to represent the codebase structure, as implemented by Aider and CodeStory Midwit Agent. In the case of Aider, the map provides a concise summary of the codebase —highlighting key classes and functions— and is included with each request sent to the LLM. [^21] Aider RepoMap: https://aider.chat/2023/10/22/repomap.html
@@
Vector stores are employed to index the codebase in submissions such as Moatless Tool, GRU, and SuperCoder. Notably, Patched.Codes Patchwork embeds the entire repository into a vector database using a Sentence-Transformers based embedding model. Aegis encodes code snippets with Cohere Embed v3 and stores them in a local FAISS vector store. In the case of Bracket.sh, the system constructs an intermediate representation of large codebases, enabling agents to operate at higher levels of layered abstraction. Other submissions, such as AppMap Navie, do not rely on embeddings or pre-built indexes but instead generate an index of the entire codebase.
@@
Other forms of representation have also been employed. For example, Amazon Q Developer Agent uses a structured XML representation of the repository’s file system, while Agentless converts the project codebase into a tree-like structure. Factory Code Droid models the codebase using a multi-resolution representation that combines explicit relationships, captured through graphs, and implicit relationships, encoded via embeddings in latent space, enabling reasoning across different levels of abstraction and a deeper understanding of the code.
@@
Nevertheless, some submissions either do not perform preprocessing or do not report it explicitly. For instance, AbanteAI MentatBot uses a Library Context tool to select relevant files, whereas Globant Code Fixer Agent, AutoCodeRover and SpecRover access the file system directly via specialized tools and APIs.

### v2

@@ 3.2.1. Human-Workflow with Fixed execution - No Agent (G1)
The entries listed in this section describe approaches that follow a predefined workflow, typically composed of two or three stages—such as localization, patch generation, and validation. These approaches execute the sequence of steps without dynamic decision-making or autonomous behavior. Moreover, none of the individual stages exhibit agentic behavior.
RAG + <model name> refers to the initial set of entries on the SWE-Bench leaderboard, introduced by its original creators as a baseline to evaluate the performance of LLMs in repairing issues from the benchmark they developed (jimenez2024SWEBenchLLMs, ). The approach follows a two-stage workflow: Localization, which uses BM25 retrieval (robertson2009probabilistic, ) to identify potentially buggy files, and Repair, where the selected LLM generates candidate patches. In addition, they fine-tuned CodeLlama (roziere2024codellama, ) on 19,000 issues to create the SWE-LLAMA model, which is used during the repair stage. Within this approach, the best results were obtained using Claude 3 Opus (4.33% on Lite, 7% on Verified).
@@
Agentless (xia2024agentlessdemystifyingllmbasedsoftware, ) is a repair workflow based on LLMs, composed of three main stages: Localization, Repair and Validation (27.33% on Lite). This workflow is similar to traditional APR, such as those implemented in the Astor framework (martinez2016Astor, ) or the state-of-the-art of traditional APR TBar (Liu2019Tbar, ), with the novelty of using LLMs to perform each task. In contrast to LLM-based Agents, in Agentless, LLMs are not used to decide how to proceed in the repair workflow. Both leaderboards also include submissions of the posterior version: Agentless-1.5 (40.67% on Lite, 50.8% on Verified).
As a pioneering approach on SWE-Bench, Agentless has been extended, modified and improved, with the results of some of these variants also submitted to the leaderboards. For example:
  • Agentless - Lite (32.3% on Lite, 42.4% on Verified), as the name suggests, is a simplification of the original approach: “Use an embedding model to retrieve relevant files from the repository (Exclusively RAG-based localization). Query the LLM to generate a repair based on the top 5 retrieved files, retrying the generation until the model outputs a valid patch.”.
@@
  • Agentless + Repograph (29.67% on Lite) corresponds to the Agentless augmented with the plugin Repograph. Repograph provides agents a structured way to analyze and interact with complex codebases, enabling detailed tracing of code dependencies, execution flow and structural relationships across the repository (ouyang2024repographenhancingaisoftware, ). Repograph information is integrated across all stages of the Agentless workflow.
@@
  • OrcaLoca + Agentless-1.5 (41% on Lite) incorporates ORCALOCA (yu2025orcaLocallmAgent, ) in the Localization stage from Agentless-1.5, while retaining the Repair, Patch Validation and Patch Selection stages. ORCALOCA is an approach focused on file and function localization. Instead of relying exclusively on LLMs to generate exploratory actions within the codebase, it introduces a dynamic action scheduling system that produces actions at varying levels of granularity and prioritizes them based on contextual relevance and urgency.
  • CodeFuse-AAIS proposes two main improvements: 1) adaptive fault localization, and 2) expansion of the sampling space by combining multiples LLMs. . Nevertheless, we could not find more details about such improvements.
@@
  • SWE-Fixer (xie2025SWEFixerTrainingopensourcellms, ) (24.67% on Lite, 32.8% on Verified) is a further simplification of Agentless, following a pipeline structure consisting of two primary stages: 1) code file retrieval, and 2) code editing. In both stages, it uses fine-tuned open-source models (Qwen 2.5) trained using Chain-of-Thought (CoT) data (wei2023chainofthought, ).
@@
Aider (26.33% on Lite) employs an iterative process composed of the following main stages: 1) File Filtering:Aider builds a repository map, a compact and expressive summary of the entire codebase, using static code analysis. This map helps the LLM understand the project structure and dependencies. The model then produces a natural language response indicating which files need to be edited and why. 2) LLM-based Code Editing:Code edits are performed using a set of carefully designed prompting strategies and configurable editing backends. 3) Linting:A static analysis phase is used to ensure that the modified code has no syntax or other fatal errors. 4) Testing:The final step consists of running all available tests to verify whether the generated patches are plausible.
@@ 3.2.2. Human-Workflow with Fixed execution - with Single Agent (G2)
Similar to the previous group (G1), the entries listed in this section describe approaches that follow a predefined workflow and execute the sequence of steps without dynamic decision-making or autonomous control. However, approaches in this group exhibit agentic behavior within one of the workflow stages.
Meta introduced the SWE-RL (wei2025SWE-RL, ) to train LLMs —resulting in the model Llama3-SWE-RL— with the goal of enabling them to autonomously recover and replicate a developer’s reasoning process. To evaluate the ability of the trained models to fix issues, the authors generated a pipeline-based scaffold named Agentless Mini with include capabilities such as file-level fault localization, generating repair edits, reproduction, test generation, and regression test selection. Agentless Mini enhances the original Agentless’ pipeline with two key improvements: 1) test generation by considering other tests, and 2) patch selection based on voting. The evaluation of this scaffold using the trained model achieved a 41.2% on SWE-Bench Verified.
@@ 3.2.3. Human-Workflow with Fixed execution - with Multiple Agents (G3)
Similar to the previous group (G2), the entries listed in this section describe approaches that follow a predefined workflow and execute the sequence of steps without dynamic decision-making or autonomous control. However, the approaches in this group exhibit varied agentic behaviors across different workflow stages, often leveraging multiple agents with complementary roles.
@@
MASAI (arora2024masaimodulararchitecturesSE, ) (27.33% on Lite, 32.6% on Verified) presents a high-level modular architecture composed of five distinct stages, including Edit Localizer, Fixer and Ranker. Each stage is executed by a dedicated LLM-powered sub-agent, which is instantiated with defined objectives and tailored strategies designed to fulfill those objectives. In particular, each sub-agent is equipped with a set of actions and a corresponding strategy, a problem-solving approach for using the LLM to address its assigned sub-task. This strategy can include methods such as vanilla completion, Chain-of-Thought(wei2023chainofthought, ) or ReAct (yao2023ReAct, ). Note that in MASAI, agentic behavior is confined to individual stages, with each sub-agent operating independently within its assigned role, rather than exercising control over the overall system workflow.
@@ 3.2.4. Human-Workflow with Scaffolded Execution - Single Agent (G4)
Similar to the previous group (G3), these approaches follow a predefined, human-authored workflow. However, instead of rigid execution, they adopt a scaffolded model where a single agent—typically an LLM—has autonomy within and across workflow stages. This allows agentic behavior in localized decisions, while the overall control flow remains fixed.
@@ 3.2.5. Human-Workflow with Scaffolded Execution - Multiple Agents (G5)
The entries of this group adopt a human-defined scaffolded workflow executed by multiple agents, where each agent operates autonomously within its assigned stage. This design enables specialization across tasks —such as localization, patch generation, and validation— while preserving a globally structured execution flow.
@@
Other approaches that report results on SWE-Bench using agentic workflows with multiple-agents, but with less publicly available detail, include Bracket.sh (53.2% on Verified) and Composio SWE-Kit (41% on Lite, 48.6% on Verified).
@@ 3.2.6. Workflow Emergent and Emergent Autonomy - Single Agent (G6)
The entries in this group operate without a predefined workflow, unlike those in the previous groups. Here, a single agent autonomously determines the execution path, selecting and ordering steps based on goals, context, and feedback.
@@
A similar approach, but using different LLMs, corresponds to the entries W&B Programmer O1 (64.6% on Verified), which was the first on SWE-Bench to evaluate the reasoning capabilities of OpenAI’s GPT-o1, and Google Jules (52.2% on Verified), which uses Google’s Gemini model.
@@ 3.2.7. Workflow Emergent and Emergent Autonomy - Multiple Agents (G7)
The entries of this group adopt an emergent workflow executed by multiple agents with predefined roles, where the coordination and execution flow are not scripted but emerge dynamically through inter-agent interaction.
@@ 3.2.8. Unclassified Submissions Due to Limited Evidence on Architectural Aspects (G8)
3.2.9. Architecture of solutions and % Resolved
@@
Architecture % Resolved % Resolved
#E Median Max #E Median Max
G1 Human-Workflow with Fixed execution - No Agent 23 27.33 46 14 24.7 50.8
G2 Human-Workflow with Fixed execution - with Single Agent - - - 4 53.7 68.2
G3 Human-Workflow with Fixed execution - with Multiple Agents 8 36.66 60.33 5 63.4 75.2
G4 Human-Workflow with Scaffolded Execution - Single Agent 16 38.16 60 10 56 70.8
G5 Human-Workflow with Scaffolded Execution - Multiple Agents 10 31.84 48.33 15 40.6 74.4
G6 Workflow Emergent and Emergent Autonomy - Single Agent 10 24.84 56.67 31 54.2 73.2
@@
G8 Unclassified Submissions 10 36.66 55 16 43.9 66.4
Table 6. Total (#E), Median and Max precision (% Resolved) of submissions grouped by the architecture.
Table 6 groups the submission according to the classification of the architectures discussed previously. For each group, it shows the total number of submissions, and the median and max precision (% Resolved).
In SWE-Bench Lite, the highest precision scores in SWE-Bench Lite were attained by submissions falling under G3 (Human-Workflow with Fixed execution - with Multiple Agents) and G4 (G4 Human-Workflow with Scaffolded Execution - Single Agent), with values of 60.33% and 60%, respectively. These two groups also stand out for the highest median precision observed among the submissions assigned to them (33.66% and 38.16%), despite similar values in other groups. This parity is confirmed by the Kruskal–Wallis test, which did not find statistically significant differences between the groups (H = 12.1876, p = 0.0579). These results indicate that well-structured, human-defined workflows, whether fixed (scripted execution) or scaffolded (allowing local autonomy), can achieve competitive performance. Moreover, the table also shows that G1 (Human-Workflow with Fixed execution - No Agent) and the mentioned G4 have the highest number of entries, with 23 and 16 submissions respectively.
In SWE-Bench Verified, the top-performing submission, classified under G3 (Human-Workflow with Fixed execution - with Multiple Agents), achieved 75.2%, though submissions from other groups recorded comparable performance levels. This group also achieved the highest overall performance, despite having a relatively small number of submissions (5). In contrast, G6 (Workflow Emergent and Emergent Autonomy – Single Agent) includes the largest number of approaches (31), with submissions reaching competitive results (up to 73.2%) and a median score that contends for second place after G3. This suggests that autonomous single-agent systems, such as those in G6, can achieve competitive repair performance, which may be attributed to the improved reasoning abilities of modern LLMs —such as Claude 4 Opus and Sonnet— that reduce the need for explicit complex scaffolding. The Kruskal–Wallis test reveals a significant difference between submitter groups (H = 19.41, p = 0.0070). Dunn’s post-hoc test shows that G1 differs significantly from G3 ($p$ = 0.0207), G4 ($p$ = 0.0125), and G6 ($p$ = 0.0003).
Answer to RQ 2 : Submissions classified under G3 (Human-Workflow with Fixed Execution – with Multiple Agents) and G4 (Human-Workflow with Scaffolded Execution – Single Agent) consistently achieved the highest performance on both SWE-Bench Lite and SWE-Bench Verified. Additionally, G6 (Workflow Emergent and Emergent Autonomy – Single Agent), which includes the largest number of submissions (31) on SWE-Bench Verified, demonstrated that autonomous single-agent approaches —empowered by the reasoning capabilities of state-of-the-art LLMs such as Claude 4 Sonnet— can also deliver competitive results without requiring complex scaffolding. Nevertheless, submissions without agents (G1) tend to show lower precision.
@@
We describe the end-to-end software maintenance pipeline proposed in Liu et al.’ survey (liu2024SurveyAgentsLLM4SE, ), which also classifies six agent-based systems that we examine in this paper: Agentless (xia2024agentlessdemystifyingllmbasedsoftware, )AutoCodeRover(Zhang2024AutoCoderRover, ), RepoUnderstander (subsequently renamed LingmaAgent(OLDma2024understandRepoUnderstander, ), MASAI (arora2024masaimodulararchitecturesSE, ), SWE-Agent(yang2024sweagentagent, ) and CodeR (xia2024dcoder, ).
@@ 3.3.1. Preprocessing
One of the initial phases commonly employed across approaches is the preprocessing of the codebase to construct a condensed representation, such as a graph or an index, that summarizes key information from the repository, including source code and, in some cases, documentation. This representation enables more efficient code navigation and supports contextual understanding. We identify different techniques used in this preprocessing phase.
@@
Several approaches evaluated on SWE-Bench adopt knowledge graph representations of the codebase. For instance, the Bytedance MarsCode Agent (liu2024marscode, ) (39.33% on Lite, 50% on Verified) analyzes and organizes the code and documentation within a repository to generate a multi-directional graph. Similarly, Nemotron-CORTEXA (sohrabizadeh2025nemotroncortexa, ) (68.2% on Verified) constructs a graph-based representation of a software repository, where nodes correspond to code files and entities —such as functions, classes, and methods— and edges capture contain and use relationships. Some approaches integrate off-the-shelf components in their solutions. For instance, Orcaloca (yu2025orcaLocallmAgent, ) (41% on Lite) employs a graph-based representation, called CodeGraph, to index code repositories and support code entity search. This graph captures hierarchical relationships —such as methods within classes— as well as reference relationships, including function calls between entities. Similarly, DARS (aggarwal2025darsdynamicactionresampling, ) (47% on Lite) utilizes RepoGraph (ouyang2024repographenhancingaisoftware, ), a hierarchical structure in which nodes represent code definitions and edges encode dependencies among them. RepoGraph is also integrated in one of Agentless’ entries (29.67% on Lite). As a last example, KGCompass (yang2025enhancing, ) (46% on Lite) incorporates a repository-aware knowledge graph designed to accurately link repository artifacts (such as issues and pull requests) with codebase entities (including files, classes, and functions). This graph is used both to narrow the search space during issue localization and to augment LLMs with relevant contextual information during patch generation. Other approaches using knowledge graphs are Alibaba Lingma Agent (33% on Lite, 28.8% on Verified) and Bytedance’s TRAE (75.2% on Verified).
@@
Repository mapping is a variant of knowledge graph, also adopted as a strategy to represent the codebase structure, as implemented by Aider (26.33% on Lite) and CodeStory Midwit Agent (62.2% on Verified). In the case of Aider, the map provides a concise summary of the codebase —highlighting key classes and functions— and is included with each request sent to the LLM. [^19] Aider RepoMap: https://aider.chat/2023/10/22/repomap.html
@@
Vector stores are employed to index the codebase in submissions such as Moatless Tool (38.33% on Lite, 70.8% on Verified), GRU (48.67% on Lite, 57% on Verified), and SuperCoder (34% on Lite). Notably, Patched.Codes Patchwork (37% on Lite) embeds the entire repository into a vector database using a Sentence-Transformers based embedding model. Aegis (30.33% on Lite) encodes code snippets with Cohere Embed v3 and stores them in a local FAISS vector store. In the case of Bracket.sh (53.2% on Verified), the system constructs an intermediate representation of large codebases, enabling agents to operate at higher levels of layered abstraction. Other submissions, such as AppMap Navie (36% on Lite, 47.2% on Verified), do not rely on embeddings or pre-built indexes but instead generate an index of the entire codebase.
@@
Some approaches utilize Repository Trees to facilitate navigation. For example, Amazon Q Developer Agent (29.67% on Lite, 65.4% on Verified) uses a structured XML representation of the repository’s file system, while most of Agentless’ entries convert the project codebase into a tree-like structure. Factory Code Droid ( 31.33% on Lite, 37% on Verified) models the codebase using a multi-resolution representation that combines explicit relationships, captured through graphs, and implicit relationships, encoded via embeddings in latent space, enabling reasoning across different levels of abstraction and a deeper understanding of the code. Repository trees are also used by SWE-Agent (56.67% on Lite, 66.6% on Verified) and Composio SWE-Kit (41% on Lite, 48.6% on Verified).
Other submissions combine different techniques. For example, Zencoder (70% on Verified) relies on Repo Grokking, a foundational engine that continuously indexes the entire codebase, capturing structural and semantic relationships through embeddings and dependency graphs. It supports advanced search capabilities—including embedding search, graph traversal, and full-text retrieval, all enhanced by LLMs for deeper contextual understanding.
@@
It must be remarked that some submissions either do not perform preprocessing or do not report it explicitly. For instance, AbanteAI MentatBot (38% on Lite) uses a Library Context tool to select relevant files, whereas Globant Code Fixer Agent (48.33% on Lite), AutoCodeRover (38.4% on Lite, 30.67% on Verified) and SpecRover (51.6% on Verified) access the file system directly via specialized tools and APIs.